Climbing ropes stretch when they catch a falling climber, thus increasing the time it takes the climber to come to rest and reducing the force on the climber. In one standardized test of ropes, an 80 kg mass falls 4.8 m before being caught by a 2.5-m-long rope.

If the net force on the mass must be kept below 11 kN, what is the minimum time for the mass to come to rest at the end of the fall?

PartA Climbing ropes stretch when they catch a falling climber, thus increasing the time it takes the climber to come to rest and reducing the force on the climber. In one standardized test of ropes, an 80 kg mass falls 4.8 m before being caught by a 2.5-m-long rope. If the net force on the mass must be kept below 11 kN, what is the minimum time for the mass to come to rest at the end of the fall? Express your answer with the appropriate units IA At0.087 SubmitPrev Previous Answers Request Answer X Incorrect; Try Again; 2 attempts remaining
Answer #1

Acceleration is given by,

a = F/m = 11000/80 = 137.5 m/s^2

Using law of conservation of energy,

Initial velocity, u = sqrt(2 x 9.8 x 4.8) = 9.7 m/s

So, time, t = 9.7/137.5 = 0.0705 sec

Comment in case any doubt please rate my answer...